If the standard deviation of the population is 25 and the sample size is 16 then the standard deviation of sampling distribution i


1
Expert's answer
2021-10-26T02:18:29-0400

Given,

"n=16\\\sigma=25"


Sampling distribution of "\bar X"



"\bar X\sim N(\mu, \dfrac{\sigma}{\sqrt n})"



Mean, "\mu = \mu_{\bar X}"


Standard deviation of sampling distribution of mean-

"\sigma_{\bar X}=\dfrac{\sigma}{\sqrt n}=\dfrac{25}{\sqrt{16}}=\dfrac{25}{4}=6.25"


Hence, the standard deviation of sampling distribution "\boxed{\sigma_{\bar X}=6.25}"